OpenStudy (anonymous):

What is the simplified expression for 2 power 2 multiplied by 2 power 3 over 2 power 4? 20 21 22 23

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[\frac{ 2^2\times2^3 }{ 2^4 }=2^2\times2^3\times 2^{-4}=2^{\left( ?+?-? \right)}=?\]

OpenStudy (anonymous):

2^2 x 2^3 x 2^-4

OpenStudy (anonymous):

2^1 right

OpenStudy (anonymous):

correct \[2^1=2\]